Circular No. 7837

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bh IN UGC 5286 M. Ganeshalingam and W. D. Li, University of California at Berkeley,

report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 24.3 (mag ∼ 17.1) and 25.3 UT (mag ∼ 17.2). SN 2002bh is located at α = 9^{h}51^{m}06^{s}.56, δ = +9^{o}00′05′′.7 (equinox 2000.0), which is 8′′.2 east and 25′′.8 south of the nucleus of UGC 5286. A KAIT image taken on Feb. 10.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002bi IN UGC 8527 Li also reports the LOTOSS discovery of another apparent supernova

in an unfiltered KAIT image taken on Feb. 25.4 UT (mag ∼ 17.1). SN 2002bi is confirmed at mag ∼ 16.2 in an earlier KAIT image taken on Jan. 31.5. The new object is located at α = 13^{h}32^{m}54^{s}.45, δ = +9^{o}31′18′′.0 (equinox 2000.0), which is 3′′.0 east and 3′′.9 north of the nucleus of UGC 8527. A KAIT image taken on 2001 May 29.3 showed nothing at this position (limiting mag ∼ 19.0).

XTE J1650–500 J. A. Tomsick and E. Kalemci, University of California at San Diego;

S. Corbel, Université Paris 7 and Laboratoire du Commissariat à l’Énergie Atomique, Saclay; and P. Kaaret, Harvard-Smithsonian Center for Astrophysics, report on a series of 30 RXTE observations of the blackholecandidate x-ray transient XTE J1650–500 (cf. IAUC 7707, 7708, 7710, 7715) taken between 2001 Dec. 27 and 2002 Feb. 24: “During this time, the 3to 20-keV light curve exhibits a large amplitude oscillation with a 13.5- to 14.0-day period about a mean flux of 6.6 × 10^{−11} erg cm^{−2} s^{−1}. The level of modulation is extremely high and roughly sinusoidal with a peak-to-peak flux variation that is close to, or possibly greater than, an order of magnitude. In addition to the four oscillation cycles we observe, the oscillation may have been present at a significantly lower level earlier in the outburst, based on the All-Sky Monitor 1.5- to 12-keV light curve. The 14-day period may be orbital, but we believe a disk precession origin (as seen in Her X-1) is more likely. Observations are encouraged to determine if the 14-day oscillation is seen at other wavelengths and to look for possible orbital or superhump signatures.”

2002 February 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green